OpenStudy (anonymous):

need help fan and medal if you help walk me through please i want to understand 2K-10 _______ =8 -7

OpenStudy (anonymous):

@malcolmmcswain @Thebossofme191 @nikeboi101 @benlindquist

OpenStudy (nikeboi101):

k in (-oo:+oo) 2*k-(10/7)-7 = 8 // - 8 2*k-(10/7)-8-7 = 0 2*k-10/7-8-7 = 0 2*k-115/7 = 0 // + 115/7 2*k = 115/7 // : 2 k = 115/7/2 k = 115/14 k = 115/14

OpenStudy (malcolmmcswain):

\[\frac{ 2k-10 }{ -7 } = 8\] What value do we need to be inserted over -7 to equal 8?

OpenStudy (malcolmmcswain):

@nikeboi101 Please don't give out direct answers. :)

OpenStudy (thebossofme191):

waznt direct

OpenStudy (malcolmmcswain):

It most certainly was.

OpenStudy (thebossofme191):

it waz step by step

OpenStudy (nikeboi101):

i techniclly ddnt i worked it out then i gav the answer

OpenStudy (anonymous):

plz dont argue

OpenStudy (anonymous):

:(

OpenStudy (malcolmmcswain):

Yes, but you don't understand how this site works. Even if you do solve it step by step, you're supposed to give help, not answers.

OpenStudy (thebossofme191):

ok sry

OpenStudy (malcolmmcswain):

Please read our code of conduct. openstudy.com/code-of-conduct

OpenStudy (nikeboi101):

ok

OpenStudy (malcolmmcswain):

It's ok. A good way to start helping someone is to ask: "What do you think?"

OpenStudy (malcolmmcswain):

Then guide them from there, and show them how to get to the answer. :)

OpenStudy (nikeboi101):

ok i will do that next time ty c:

OpenStudy (anonymous):

i think it is -7??

OpenStudy (malcolmmcswain):

Thanks. c:

OpenStudy (anonymous):

i dont understand

OpenStudy (nikeboi101):

oh

OpenStudy (anonymous):

math hates me XD

OpenStudy (thebossofme191):

same

OpenStudy (anonymous):

lol

OpenStudy (thebossofme191):

but u already know that

OpenStudy (thebossofme191):

lol

OpenStudy (nikeboi101):

ok ima do it step by step ok?

OpenStudy (anonymous):

@isherron

OpenStudy (anonymous):

opps

OpenStudy (nikeboi101):

allright Step 1 : 10 Simplify —— -7

OpenStudy (nikeboi101):

10 (2k - ——) + 8 -7

OpenStudy (nikeboi101):

Step 2 : Rewriting the whole as an Equivalent Fraction :

OpenStudy (anonymous):

k

OpenStudy (nikeboi101):

Rewrite the whole as a fraction using -7 as the denominator : 2k 2k • -7 2k = —— = ——————— 1 -7

OpenStudy (nikeboi101):

2k • 7 - (10 • -1) 14k + 10 —————————————————— = ———————— 7 7 Equation at the end of step 2 : (14k + 10) —————————— + 8 7

OpenStudy (thebossofme191):

ayy im learning too lol

OpenStudy (anonymous):



OpenStudy (nikeboi101):

Step 3 : Rewriting the whole as an Equivalent Fraction : 3.1 Adding a whole to a fraction Rewrite the whole as a fraction using 7 as the denominator :

OpenStudy (nikeboi101):

8 8 • 7 8 = — = ————— 1 7

OpenStudy (nikeboi101):

Step 4 : Pulling out like terms :

OpenStudy (nikeboi101):

14k + 10 = 2 • (7k + 5)

OpenStudy (anonymous):

10,-7

OpenStudy (nikeboi101):

Adding fractions that have a common denominator : 4.2 Adding up the two equivalent fractions 2 • (7k+5) + 8 • 7 14k + 66 —————————————————— = ———————— 7 7

OpenStudy (anonymous):

are the like terms right?

OpenStudy (nikeboi101):

Step 5 : Pulling out like terms : 5.1 Pull out like factors : 14k + 66 = 2 • (7k + 33) Final result : 2 • (7k + 33) ————————————— 7

OpenStudy (nikeboi101):

yea

OpenStudy (anonymous):

k XD

OpenStudy (nikeboi101):

and if you still dont understand go over these steps and you will eventually get it c:

OpenStudy (anonymous):

ok thank you so much XD

OpenStudy (nikeboi101):

YOUR WELCOME

OpenStudy (anonymous):

XD